one internet cafe charges $1 for 15 min and $0.20 per page for printing. Another cafe charges $2.00 per hour and $0.25 per page for printing. If you wanted to use the cafe for 1 hour, how many pages would you need to print in order for the cafe prices to be equal?
Answer by mananth(16946)

internet cafe I charges $1 for 15 min
So $4 per hour
$0.20 per page for printing.
Let number of pages printed be x
C I = 4+0.2x
cafe II charges $2.00 per hour and $0.25 per page for printing.
C II = 4 +0.25x
for the cafe prices to be equal
4+0.2x=2+0.25x
2= 0.05x
x= 40 pages